OpenStudy (anonymous):

Julissa is running a 10-kilometer race at a constant pace. After running for 18 minutes, she completes 2 kilometers. After running for 54 minutes, she completes 6 kilometers. Her trainer writes an equation letting t, the time in minutes, represent the independent variable and k, the number of kilometers, represent the dependent variable.

OpenStudy (anonymous):

So you're supposed to write the equation?

OpenStudy (anonymous):

Points are (independent variable, dependent variable). For this problem that's (minutes, kilometers). So what are the two points you're given?

OpenStudy (anonymous):

(18,2) (54,6)

OpenStudy (anonymous):

yes. Do you know how to find the equation of a line through two points?

OpenStudy (anonymous):

i'm not completely sure. can you walk me through it?

OpenStudy (anonymous):

Start by finding slope \[m=\frac{ y _{2} -y _{1}}{ x _{2} -x _{1} }\]

OpenStudy (anonymous):

1/9 = slope

OpenStudy (anonymous):

yes. so the normal equation of a line is y = mx + b, but they switched you letters to k and t so your equation is k = mt + b.

OpenStudy (anonymous):

Plug in m and one of the points for k and t and solve for b

OpenStudy (anonymous):

k-2=1/9(t-18) ?

OpenStudy (anonymous):

yes